1. A 4 letter word is written down by taking letters from the word KANYAKUMARI. What is the chance of the word to be of all different letters?


2. A card is drawn from a pack of 52 cards. The card is then replaced and the pack is reshuffled. It is done 6 times. Find the chance that the cards drawn are 2 hearts, 2 diamonds & 2 blacks.

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Abhipsa

KANYAKUMARI =2 K +3 A + 1N + 1Y + 1U + 1M + 1R +1I

Case letter does not include A and K

total number of 4 letter word can be formed =6C4  4!................1

total number of 4 letter word that has different letter can be formed =6C4  4!................2

 

case letter include A but not K

 

total number of 4 letter word can be formed =1*6C3  4! + 1*1*6C2  4!/2!   +1*1*1*6C1  4!/3! ( when 1 A ,2A and 3A are included respestively)...............3

total number of 4 letter word that has different letter can be formed =1*  6C3  4! ...........4


case letter include K but not A

 

total number of 4 letter word can be formed =1*6C3  4! + 1*1*6C2  4!/2!   + ( when 1 K ,2K  are included respestively).................5

total number of 4 letter word that has different letter can be formed =1*  6C3  4!...............6

case letter  K and  A both included

 

total number of 4 letter word can be formed =1*1*6C2  4! + 1*2*6C1  4!/2!   + 1*3*6C0  4!/3! +2*1*6C2  4!/2! +2*2*6C0  4!/2! 2!( when 1 K 1A ,1K2A,1K3A,2k1A,2K 2A   are included respestively).................7

total number of 4 letter word that has different letter can be formed =1*1*  6C2  4!...............8

 

so pobability =(2+4+6+8)/(1+3+5+7)

   put value of above equation and calculate


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

Abhipsa Mishra
14 years ago
Thank you for the answer. Please answer the second question also.
Abhipsa Mishra
14 years ago

Thank you for your answer. Please answer the second question that i've posted.